Question 19 1 pts If the government were to impose a price celling of s6, which statement is correct sis ZS12 $6 $3 10 20 30 50 QUANTITY 0 There would be a shortage of 20 units and a deadweight loss of 30 O There would be a surplus of 20 units and Consumer Surplus will be smaller than before the price ceiling The price ceiling in not binding therefore deadweight loss is equal to 0 O There would be a shortage of 20 units and Producer Surplus will be greater than before the price ceiling Question 20

if government were to impose a price ceiling or 6, which statement is correct
Answer #1

Answer
The market is in equilibrium at Qd=Qs=30 and P=$9
the Pe>Pc
so the price ceiling is effective e and decreases market price and increases demand but the supply decreases so there is a shortage
shortage=Qd-Qs=40-20=20 units
Dead weight loss=0.5*(12-6)*(30-20)=30

the Producer surplus decreases after price ceiling.
Option first.